# Origins of the UV continuum and Balmer emission lines in Little Red Dots: observational validation of dense gas envelope models enshrouding the AGN

### v1

@@ IV.2 Future tests of envelope models
Although the blackbody envelope model is currently the most preferred, other types of envelope models are not fully ruled out, and these models can be further validated with future observations. One of the promising approaches is to explore the time variability of LRDs (e.g., Furtak et al., 2025; Kokubo and Harikane, 2025; Ji et al., 2025). Since emissions powered by AGN accretion can have time variability while stellar-origin emissions are not or weakly variable, we can distinguish which emission components in LRDs originate from AGNs or stars with their variability. For example, the blackbody envelope model predicts time variability only in the rest optical continuum but not in the emission line luminosities, while the cocoon model predicts both in the rest optical continuum and the broad line luminosity. Recently, Zhang et al. (2025) utilized the time delay effect between multiple images of one single LRD due to the strong gravitational lensing and reported long-term ( in the observer frame) variability in the photometric colors, yet it is not clear whether the broad line component also shows the variability. Deep spectroscopic observations of multiply-lensed LRDs will thus be an important diagnostic to validate the envelope models.
Another key observations would be to probe the structures of BLRs in LRDs. Normal AGNs in the local universe are known to have a stratified ionized structure in BLRs, where higher ionization potential broad lines reside in the inner radii (e.g., Peterson, 2006; Gaskell, 2009; Kollatschny and Zetzl, 2013). A few previous studies on LRDs found that broad low ionization-potential lines have lower FWHMs than broad Balmer emission lines (e.g., O i 8448 or Fe ii; Kokorev et al. 2025; Lin et al. 2025; Tripodi et al. 2025), suggesting that BLRs in LRDs could also have similar stratified structures. This can be naturally explained by the cocoon envelope model or the non-spherical envelope model, while an extra explanation is needed in the blackbody envelope model. However, the stratified BLR structures in LRDs have been observed only in lines with ionization potentials lower than Balmer lines that can also be emitted by Ly/Ly fluorescence process, which makes the detailed ionization structures in LRD BLRs remain unclear. In particular, finding high ionization-potential lines such as C iv1550 with broader line widths than Balmer lines in LRDs would place a severe constraint against the blackbody envelope model.
@@ IV.3 Duty cycles of LRDs
Our analysis indicates that the UV continuum observed in LRD spectra originate from the stellar emissions. If this is the case, the UV luminosity function (LF) of LRDs directly reflects that of their stellar component. Previous studies with photometrically selected LRDs (Kokorev et al., 2024; Kocevski et al., 2025) showed that the UV LF of LRDs at is approximately two orders of magnitude lower in normalization than that of Lyman-break galaxies at the same redshifts, while the LF shapes are remarkably similar across a wide magnitude range (from to mag; e.g., Bouwens et al., 2021; Harikane et al., 2022).
The similar LF shapes but different normalizations suggest that only of rest-UV-selected star-forming galaxies at appear as LRD host galaxies. This fraction can be interpreted as the duty cycle of LRDs, yielding at these redshifts. Such a duty cycle corresponds to a typical LRD lifetime of of the Hubble time, or Myr. The estimate agrees well with an independent constraint on based on the characteristic timescale of stellar feedback, which is expected to quench gas supply to the nuclei and thus disperse the thick envelope structure (Inayoshi et al., 2025a).
@@ IV.4 BH mass estimator for LRDs
In the blackbody envelope model (Figure 1a), even the broad Balmer emission lines are not primarily powered by the AGN accretion disk, but instead by young massive stars in the compact stellar cluster. As a result, no direct emission from the accretion disk is observable, and single-epoch virial BH mass estimates based on the - relation calibrated by reverberation mapping of nearby AGNs become inappropriate (e.g., Greene and Ho, 2005; Bentz et al., 2013). Instead, the broad-line clouds illuminated by stellar radiation reside near the surface of the gaseous envelope and are gravitationally bounded by the central BH. Thus, the characteristic radius of the BLR can be approximated by the photospheric radius of the envelope , allowing the BH mass to be inferred from and the line width of the broad H emission line.
Figure 5 compares the BH mass measurements of our LRD sample derived from the standard - relation and those inferred from the blackbody photospheric radius. In the former case, we estimate BH masses following the prescription of Reines et al. (2013), which uses the broad H luminosity as a proxy for under the assumption of the standard AGN – relation (the -axis of Figure 5). In the latter case, we measure the blackbody temperature () and the total envelope luminosity () by fitting blackbody spectra to the Prism spectra of the sample LRDs, and estimate using the Stefan-Boltzmann law. Assuming that the broad-line clouds are virialized, we then compute the corresponding BH with (the -axis of Figure 5). Plots are color-coded by the best-fit blackbody temperature. It shows that the BB temperatures are tightly clustered around K independently of the BH mass, which is in line with the prediction by the blackbody envelope model (e.g., Kido et al., 2025; de Graaff et al., 2025b; Umeda et al., 2025).
The comparison demonstrates that BH masses inferred under the blackbody envelope model are systematically lowered by one order of magnitude than those obtained using the conventional single-epoch method. This reduction relieves the “overmassive” nature of BHs in LRDs relative to the local BH-host mass correlations (e.g., Furtak et al., 2024b; Maiolino et al., 2024; Kocevski et al., 2025). We highlight that this decrease rises from the fundamentally different BLR size estimates, rather than from additional line broadening due to electron or resonant scattering (e.g., Rusakov et al., 2025a). It is thus possible that the BH mass may be even smaller by a further dex, if the additional line broadening due to scattering takes place in the blackbody envelope model.
@@ V Summary
Given these observational results, the blackbody envelope model (e.g., Inayoshi et al., 2025a) is the most favored, where the BH is enshrouded by a completely opaque gaseous envelope, and all UV, narrow, and broad lines are predominantly powered by young massive stars in the compact nuclear region. The cocoon envelope model (e.g., Naidu et al., 2025), which predicts the AGN-powered broad lines escaping through the dense gas cocoon, is not completely rejected, although it requires several non-trivial and fine-tuned assumptions to explain the full set of observations (Section IV.1).
@@
Given the continuous distribution of LRDs and LBDs in the - diagram, LRDs are not a distinct population, but rather they could gradually evolve to normal blue unobscured AGNs. Non-spherical envelope models (e.g., Lin et al., 2025) may play an important role during this transition phase from LRDs to normal AGNs and provide a physical link to the unified model of typical AGNs (e.g., see Netzer, 2015).

### v2

@@ IV.2 Future tests of envelope models
Although the blackbody envelope model is currently the most preferred, other types of envelope models are not fully ruled out, and these models can be further validated with future observations. One of the promising approaches is to explore the time variability of LRDs (e.g., L. J. Furtak et al., 2025; M. Kokubo & Y. Harikane, 2025; X. Ji et al., 2025). Since emissions powered by AGN accretion can have time variability while stellar-origin emissions are not or weakly variable, we can distinguish which emission components in LRDs originate from AGNs or stars with their variability. For example, the blackbody envelope model predicts time variability only in the rest optical continuum but not in the emission line luminosities, while the cocoon model predicts both in the rest optical continuum and the broad line luminosity. Recently, Z. Zhang et al. (2025) utilized the time delay effect between multiple images of one single LRD due to the strong gravitational lensing and reported long-term ( in the observer frame) variability in the photometric colors, yet it is not clear whether the broad line component also shows the variability. Deep spectroscopic observations of multiply-lensed LRDs will thus be an important diagnostic to validate the envelope models.
Another key observations would be to probe the structures of BLRs in LRDs. Normal AGNs in the local universe are known to have a stratified ionized structure in BLRs, where higher ionization potential broad lines reside in the inner radii (e.g., B. M. Peterson, 2006; C. M. Gaskell, 2009; W. Kollatschny & M. Zetzl, 2013). A few previous studies on LRDs found that broad low ionization-potential lines have lower FWHMs than broad Balmer emission lines (e.g., O i 8448 or Fe ii; V. Kokorev et al. 2025; X. Lin et al. 2025; R. Tripodi et al. 2025), suggesting that BLRs in LRDs could also have similar stratified structures. This can be naturally explained by the cocoon envelope model or the non-spherical envelope model, while an extra explanation is needed in the blackbody envelope model. However, the stratified BLR structures in LRDs have been observed only in lines with ionization potentials lower than Balmer lines that can also be emitted by Ly/Ly fluorescence process, which makes the detailed ionization structures in LRD BLRs remain unclear. In particular, finding high ionization-potential lines such as C iv1550 with broader line widths than Balmer lines in LRDs would place a severe constraint against the blackbody envelope model.
@@ IV.3 Duty cycles of LRDs
Our analysis indicates that the UV continuum observed in LRD spectra originate from the stellar emissions. If this is the case, the UV luminosity function (LF) of LRDs directly reflects that of their stellar component. Previous studies with photometrically selected LRDs (V. Kokorev et al., 2024; D. D. Kocevski et al., 2025) showed that the UV LF of LRDs at is approximately two orders of magnitude lower in normalization than that of Lyman-break galaxies at the same redshifts, while the LF shapes are remarkably similar across a wide magnitude range (from to mag; e.g., R. J. Bouwens et al., 2021; Y. Harikane et al., 2022).
The similar LF shapes but different normalizations suggest that only of rest-UV-selected star-forming galaxies at appear as LRD host galaxies. This fraction can be interpreted as the duty cycle of LRDs, yielding at these redshifts. Such a duty cycle corresponds to a typical LRD lifetime of of the Hubble time, or Myr. The estimate agrees well with an independent constraint on based on the characteristic timescale of stellar feedback, which is expected to quench gas supply to the nuclei and thus disperse the thick envelope structure (K. Inayoshi et al., 2025a).
@@ IV.4 BH mass estimator for LRDs
In the blackbody envelope model (Figure 1a), even the broad Balmer emission lines are not primarily powered by the AGN accretion disk, but instead by young massive stars in the compact stellar cluster. As a result, no direct emission from the accretion disk is observable, and single-epoch virial BH mass estimates based on the - relation calibrated by reverberation mapping of nearby AGNs become inappropriate (e.g., J. E. Greene & L. C. Ho, 2005; M. C. Bentz et al., 2013). Instead, the broad-line clouds illuminated by stellar radiation reside near the surface of the gaseous envelope and are gravitationally bounded by the central BH. Thus, the characteristic radius of the BLR can be approximated by the photospheric radius of the envelope , allowing the BH mass to be inferred from and the line width of the broad H emission line.
Figure 5 compares the BH mass measurements of our LRD sample derived from the standard - relation and those inferred from the blackbody photospheric radius. In the former case, we estimate BH masses following the prescription of A. E. Reines et al. (2013), which uses the broad H luminosity as a proxy for under the assumption of the standard AGN – relation (the -axis of Figure 5). In the latter case, we measure the blackbody temperature () and the total envelope luminosity () by fitting blackbody spectra to the Prism spectra of the sample LRDs, and estimate using the Stefan-Boltzmann law. Assuming that the broad-line clouds are virialized, we then compute the corresponding BH with (the -axis of Figure 5). Plots are color-coded by the best-fit blackbody temperature. It shows that the BB temperatures are tightly clustered around K independently of the BH mass, which is in line with the prediction by the blackbody envelope model (e.g., D. Kido et al., 2025; A. de Graaff et al., 2025c; H. Umeda et al., 2025).
The comparison demonstrates that BH masses inferred under the blackbody envelope model are systematically lowered by one order of magnitude than those obtained using the conventional single-epoch method. This reduction relieves the “overmassive” nature of BHs in LRDs relative to the local BH-host mass correlations (e.g., L. J. Furtak et al., 2024a; R. Maiolino et al., 2024; D. D. Kocevski et al., 2025). We highlight that this decrease rises from the fundamentally different BLR size estimates, rather than from additional line broadening due to electron or resonant scattering (e.g., V. Rusakov et al., 2025b). It is thus possible that the BH mass may be even smaller by a further dex, if the additional line broadening due to scattering takes place in the blackbody envelope model.
@@ V Summary
Given these observational results, the blackbody envelope model (e.g., K. Inayoshi et al., 2025a) is the most favored, where the BH is enshrouded by a completely opaque gaseous envelope, and all UV, narrow, and broad lines are predominantly powered by young massive stars in the compact nuclear region. The cocoon envelope model (e.g., R. P. Naidu et al., 2025), which predicts the AGN-powered broad lines escaping through the dense gas cocoon, is not completely rejected, although it requires several non-trivial and fine-tuned assumptions to explain the full set of observations (Section IV.1).
@@
Given the continuous distribution of LRDs and LBDs in the - diagram, LRDs are not a distinct population, but rather they could gradually evolve to normal blue unobscured AGNs. Non-spherical envelope models (e.g., X. Lin et al., 2025) may play an important role during this transition phase from LRDs to normal AGNs and provide a physical link to the unified model of typical AGNs (e.g., see H. Netzer, 2015).